# Conditional Density Estimation with Neural Networks for Bias Correction of Multivariate Climate Model Data

### v1

@@ Abstract
Global Climate Models (GCMs) are numerical models that simulate complex physical processes within the Earth’s climate system, and are essential for understanding and predicting climate change. However, GCMs suffer from systemic biases due to assumptions about and simplifications made to the underlying physical processes. GCM output therefore needs to be bias corrected before it can be used for future climate projections. Most common bias correction methods, however, cannot preserve spatial, temporal, or inter-variable dependencies. We propose a new bias correction method based on conditional density estimation for the simultaneous bias correction of daily precipitation and maximum temperature data obtained from gridded GCM spatial fields. The Vecchia approximation is employed to preserve dependencies in the data, and conditional density estimation is carried out using semi-parametric quantile regression. Illustration on historical data from 1951–2014 over two $5\times 5$ spatial grids in the US indicate that our method can preserve key marginal and joint distribution properties of precipitation and maximum temperature, and predictions obtained using our approach are better calibrated compared to predictions using asynchronous quantile mapping and canonical correlation analysis, two commonly used alternative bias correction approaches.
Key words: Bias correction, Climate modeling, Geostatistics, Neural networks, Vecchia approximation.
@@ 1 Introduction
Global Climate Models (GCMs) are numerical models to explain complex interactions between the ocean, atmosphere and land surface. GCMs are used to study long-term climate trends and to provide climate projections based on possible future scenarios associated with, e.g., greenhouse gas and aerosol emissions, land use, and population growth [^1] https://www.climate.gov/maps-data/climate-data-primer/predicting-climate/climate-models. The latest generation of GCMs, collectively known as the Climate Model Intercomparison Project version 6 (CMIP6) (Eyring et al., 2016), contains the output from coordinated experiments that incorporate a range of possible scenarios. GCMs are used for both surface water and groundwater availability over the planning period (Singh et al., 2015; Seo et al., 2016). To develop these assessments of future water availability, GCM projections, particularly for precipitation and temperature, are forced through hydrologic models, which are subsequently used as an input into a reservoir model (Singh et al., 2015; Seo et al., 2016). Projected inflow into reservoirs are typically used for allocating water for multipurpose uses that includes municipal, irrigation, hydropower and flood control. GCM projections for other variables, particularly temperature and humidity, are also critical for quantifying energy demand (Eshraghi et al., 2021).
@@
Recent computational advances and parameterization of physical processes (e.g., cloud physics) have improved the resolution and accuracy of GCM projections under potential emission scenarios (Ramirez-Villegas et al., 2013; Strandberg and Lind, 2021). Despite their usefulness, GCM outputs have some key drawbacks which need to be addressed before they can be used in climate change studies. First, they have systematic biases due to model errors that arise from assumptions about and simplifications made to the underlying physical processes (Stevens and Bony, 2013; Du et al., 2022). These biases lead to discrepancies between observed and model data, and an inability to estimate the observed statistics of climate variables (Bhowmik and Sankarasubramanian, 2019). Such discrepancies between modeled and observed data result in error propagation in sectoral analysis such as water management (Singh et al., 2015; Seo et al., 2016). In addition, hydrologic projections developed using GCM projections may exhibit uncertainty due to hydrologic models (Bhowmik et al., 2017; Seo et al., 2019) or from (long-term) scenarios (Lettenmaier et al., 1999; Hawkins and Sutton, 2011), adding challenges in sectoral application. Second, the coarse resolution of GCMs often fails to fully resolve the effect of topography on hydroclimatic variables, especially precipitation. While some CMIP6 models have a 50 km spatial resolution, most are between 100–250 km. At these resolutions, topography is poorly represented, and GCMs are therefore unable to accurately capture the atmospheric processes that drive precipitation and precipitation extremes (Li et al., 2011; O’Brien et al., 2016). These two problems are addressed by bias-correcting and downscaling the GCM output, respectively, carried out over the historical period where both observational and model data are available. The bias correction and downscaling methods can then be used to obtain fine-scale, unbiased projections of future climate.
@@
The remainder of this manuscript focuses on the bias correction of daily precipitation (PRCP) and maximum temperature (TMAX) data based on CMIP6 GCMs. Traditional methods like quantile mapping (QM) have been used to adjust the distribution of modeled data and align it with observed data by mapping quantiles from the model to corresponding quantiles in the observational record (Stoner et al., 2013; Thrasher et al., 2012; Maraun, 2013). While effective for single-variable adjustments, QM often fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). Multivariate bias correction methods aim to preserve the relationships between climatic variables. While multivariate QMs attempt to address this issue (Piani and Haerter, 2012; Cannon, 2018), QM is inappropriate for the bias correction of future projections as it distorts future climate change in a nonphysical way (e.g., Maurer and Pierce, 2014). Other commonly used methods such as bias corrected spatial downscaling (BCSD, Wood et al., 2002), bias corrected constructed analogues (BCCA, Maurer et al., 2010), and multivariate adapted constructed analogues (MACA, Abatzoglou and Brown, 2012) also employ QM. Multivariate methods like localized constructed analogs (LOCA; Pierce et al., 2014, 2023) and asynchronous canonical correlation analysis (CCA; Bhowmik et al., 2017) offer improvements by better capturing the temporal covariability in climate variables (e.g., cross-correlation), but do not explicitly consider spatial dependence during bias correction. Finally, recent deep learning-based methods, like DeepSD (Vandal et al., 2017), leverage the power of deep learning algorithms to model complex, nonlinear relationships between climate variables and local physiographical information such as elevation, but have not been evaluated for spatio-temporal dependency in multiple climate variables. Additionally, none of the aforementioned methods, with the exception of LOCA v2 (Pierce et al., 2023), are geared towards reproducing observed extreme quantiles of climate variables. This limits their practical utility for studying climate extremes.
@@
The remainder of this paper is organized as follows. Section 2 introduces the data and methods for our study. Section 3 contains a case study for bias correction of climate model data at two different regions of the US. Section 4 concludes.
@@ 2.1 Overview
Section 2.2 describes the different datasets used in our study, including resolution and pre-processing used to align the different datasets on a common spatial grid. Section 2.3 introduces the SPCDE approach for bias correction of TMAX and PRCP at a single grid cell. Sections 2.4 and 2.5 extend SPCDE to spatial and spatiotemporal settings, respectively. Section 2.6 outlines SPQR, the conditional density estimator that is used for SPCDE. Finally, Section 2.7 describes the metrics used to assess the quality of the bias-corrected data.
@@ 2.2 Temperature and precipitation datasets
2.3 Bias correction model for data at a single location
We begin by stating our model for the general multivariate case in this section; explicit formulations for spatial and spatiotemporal data are described in subsequent sections. We build a statistical model for $p$ continuous responses, $\mbox{\bf Y}=(Y_{1},...,Y_{p})$ at a single location, which are assumed to have invertible marginal quantile and distribution functions that depends on variables Z (e.g., season or elevation). The model is described simultaneously for both climate model output and observational data, with an auxiliary variable $X=0$ indicating the observation is from the climate model and $X=1$ indicating it is from a station measurement.
@@ 2.3 Bias correction model for data at a single location
While (4) holds for any ordering of the $p$ variables, orderings which have a physical interpretation are preferable. For example, in the simple case of TMAX and PRCP at a single location (i.e., $p=2$), a natural ordering would be based on choosing $Y_{1}$ as TMAX and $Y_{2}$ as PRCP. This decomposes the conditional distribution of PRCP and TMAX as a conditional and a marginal distribution, assuming PRCP depends on TMAX. Physically, TMAX and PRCP are coupled, and choosing $Y_{1}$ as PRCP and $Y_{2}$ as TMAX also provides a valid decomposition. However, we assume that PRCP depends on TMAX for somewhat practical reasons. Intuitively, the distribution of TMAX has a much simpler functional form compared to PRCP. PRCP has a preponderance of zeros, and is therefore harder to estimate compared to TMAX. Our proposed dependence structure mitigates this by using TMAX as a predictor for PRCP. Our formulation is also validated in practice; SPCDE where TMAX depended on PRCP led to noticeably worse predictions over both the training and testing periods.
@@
Algorithm 1 Bias correction of TMAX and PRCP at a single location
SPQR estimates of $F_{1},Q_{1},F_{2},Q_{2}$
Data $\mbox{\bf Y},X$
procedure TMAX bias correction
@@
procedure PRCP Bias correction
@@ 2.4 Extension to spatial data
Consider our problem of interest - bias correction of TMAX and PRCP at multiple locations. If there are $n$ locations, a reasonable approach would be to first order them (e.g., based on their geographical coordinates), and then order TMAX and PRCP at each location. This could be represented by $\mbox{\bf Y}=(Y_{1},\ldots,Y_{p})$ for $p=2n$, and theoretically, Algorithm 1 from Section 2.3 can be applied. However, as the number of locations increase, so does the number of features $\mbox{\bf Y}_{(j)}$. This soon leads to computational problems, and directly applying (1) to large spatial (and spatiotemporal) data is computationally prohibitive. To mitigate this, we approximate (1) by means of a Vecchia approximation (Vecchia, 1988; Stein et al., 2004; Datta et al., 2016; Katzfuss and Guinness, 2021):
@@ 2.5 Extension to spatiotemporal data
In this study, we assume the conditional quantile function to be first order Markov in time, i.e., $T^{\prime}=1$. This implies that the Vecchia neighbor set for $Y_{1lt}$ is $\mbox{\bf Y}_{(1lt)}=\{Y_{1l,t-1},Y_{1,l-1,t},...,Y_{11t}\}$, and for $Y_{2lt}$ is $\mbox{\bf Y}_{(2lt)}=\{Y_{2l,t-1},Y_{1l,t-1},Y_{1,l,t},\ldots,Y_{1,l-m,t},Y_{2,l-1,t},...,Y_{2,l-m,t}\}$. These are straightforward extensions of the Vecchia neighbor sets presented in Section 2.4, with TMAX on day $t$ being additionally depending on TMAX on day $t-1$, and PRCP on day $t$ additionally depending on both PRCP and TMAX from day $t-1$. The conditional QF $Q$, CDF $F$, and calibration function $C$, for $\mbox{\bf Y}_{lt}$ are defined identically to Section 2.3; these do not depend on $t$ due to the stationarity assumption. Algorithm 2 details the calibration methodology for spatiotemporal data.
Algorithm 2 Spatiotemporal bias correction for TMAX and PRCP
SPQR estimates of $F_{1l},Q_{1l},F_{2l},Q_{2l}$, $l\in\{1,\ldots,n\}$
@@
     procedure PRCP Bias correction
@@ 2.6 Conditional density estimation using SPQR
$\Pi_{K}(\mbox{\bf X}):=\mbox{\bf x}^{(H+1)}$ $=\mbox{softmax}(\mbox{\bf W}^{(H+1)}\mbox{\bf x}^{(H)}+\mbox{\bf b}^{(H+1)})$ (8)
@@
$\frac{\partial Q(\tau|\phi,\mbox{\bf X})}{\partial\tau}>0,\ \forall\ \mbox{\bf X}$
and does not require any second-stage monotonization treatment Bondell et al. (2010).
Our approach connects CDE (Hyndman et al., 1996; Li and Racine, 2006) and bias correction in a unique manner, and in the process mitigates several drawbacks of commonly used bias correction methods. A key aspect of the proposed SPCDE approach (discussed in further detail in Sections 2.32.5) is the estimation of the joint distribution of GCM and observational data within a single SPQR model, and thereafter using cumulative density function (CDF) and quantile function (QF) transforms to bias correct the model data. Similar CDF based approaches have been used for bias correcting future climate projections (e.g., Michelangeli et al., 2009; Li et al., 2010); however, these approaches estimate the CDF of the GCM and observed data independently of each other. Since SPCDE models the joint tail distribution of observed and GCM data, we anticipate that it will provide more realistic bias-correction of extremes. There are also similarities between SPCDE and normalizing flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021; Kobyzev et al., 2021), which uses a series of invertible transformations (known as flows) to learn a mapping from simple univariate distributions to more complex joint distributions. Normalizing flows are themselves a form of CDE. Conversely, SPCDE can be considered a normalizing flow where the flows are a sequence of CDFs and QFs. However, normalizing flows do not consider an explicit spatial dependence structure, whereas it is straightforward to specify in our proposed approach.

### v2

# Spatiotemporal Density Correction of Multivariate Global Climate Model Projections using Deep Learning

@@ Abstract
Global Climate Models (GCMs) are numerical models that simulate complex physical processes within the Earth’s climate system and are essential for understanding and predicting climate change. However, GCMs suffer from systemic biases due to simplifications made to the underlying physical processes. GCM output therefore needs to be bias corrected before it can be used for future climate projections. Most common bias correction methods, however, cannot preserve spatial, temporal, or inter-variable dependencies. We propose a new semi-parametric conditional density estimation (SPCDE) for density correction of the joint distribution of daily precipitation and maximum temperature data obtained from gridded GCM spatial fields. The Vecchia approximation is employed to preserve dependencies in the observed field during the density correction process, which is carried out using semi-parametric quantile regression. The ability to calibrate joint distributions of GCM projections has potential advantages not only in estimating extremes, but also in better estimating compound hazards, like heat waves and drought, under potential climate change. Illustration on historical data from 1951–2014 over two $5\times 5$ spatial grids in the US indicate that SPCDE can preserve key marginal and joint distribution properties of precipitation and maximum temperature, and predictions obtained using SPCDE are better calibrated compared to predictions using asynchronous quantile mapping and canonical correlation analysis, two commonly used bias correction approaches.
Key words: Bias correction, Climate modeling, Density correction, Geostatistics, Neural networks, Vecchia approximation.
@@ 1 Introduction
Global Climate Models (GCMs) are numerical models to explain complex interactions between the ocean, atmosphere, and land surface. GCMs are used to study long-term climate trends and to provide projections based on possible future scenarios associated with, e.g., greenhouse gas and aerosol emissions, land use, and population growth [^1] https://www.climate.gov/maps-data/climate-data-primer/predicting-climate/climate-models. The latest generation of GCMs are collectively known as the Climate Model Intercomparison Project version 6 (CMIP6; Eyring et al., 2016). It contains the outputs, primarily ocean, atmospheric, and land surface conditions, as ensembles at a coarser spatial scale for further analyzing impacts on various sectors. For instance, GCM forcings, typically precipitation and temperature, are used to assess for both surface water and groundwater availability over the planning period (Singh et al., 2015; Seo et al., 2016). To develop these assessments of future water availability, GCM projections, particularly for precipitation and temperature, are forced through hydrologic models, which are subsequently used as an input into a reservoir model (Singh et al., 2015; Seo et al., 2016). Projected inflow into reservoirs are typically used for allocating water for multipurpose uses that includes municipal, irrigation, hydropower and flood control. GCM projections for other variables, particularly temperature and humidity, are also critical for quantifying energy demand (Eshraghi et al., 2021).
@@
Recent computational advances and parameterization of physical processes (e.g., cloud physics) have improved the resolution and accuracy of GCM projections under potential emission scenarios (Ramirez-Villegas et al., 2013; Strandberg and Lind, 2021). Despite these advances, GCM projections have some key drawbacks which need to be addressed before they can be used in climate change impact studies. First, they have systematic biases due to model errors that arise from simplifications made to the underlying physical processes (Stevens and Bony, 2013; Du et al., 2022). These biases lead to discrepancies between observed and model data resulting in an inability to estimate the observed statistics of climate variables (Bhowmik and Sankarasubramanian, 2019). Such discrepancies between modeled and observed data also result in error propagation in sectoral analysis (e.g., water management) (Singh et al., 2015; Seo et al., 2016). In addition, hydrologic projections developed using GCM projections may exhibit uncertainty due to these imprecise forcings (Bhowmik et al., 2017; Seo et al., 2019) or from (long-term) scenarios (Lettenmaier et al., 1999; Hawkins and Sutton, 2011), adding challenges in sectoral application. Second, the coarse resolution of GCMs often fails to fully resolve the effect of topography or sub-grid variability on hydroclimatic variables, especially in projecting precipitation. While some CMIP6 models have a 50 km spatial resolution, most are between 100–250 km. At these resolutions, topography, land use, and land cover are poorly represented; GCMs are therefore unable to accurately capture the land surface and atmospheric feedback that drive local hydroclimatic extremes (Li et al., 2011; O’Brien et al., 2016). These two problems are addressed by bias-correcting and downscaling the GCM output, respectively, carried out over the historical period where both observational and model data are available. The bias correction and downscaling models can then be used to obtain fine-scale, unbiased projections of future climate.
@@
Traditional bias correction methods like quantile mapping (QM) have been used to adjust the distribution of GCM data by mapping quantiles from the GCM to the corresponding quantiles in the observational record (Stoner et al., 2013; Thrasher et al., 2012; Maraun, 2013). While effective for single-variable adjustments, QM often fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). Multivariate bias correction methods aim to preserve the relationships between climatic variables (Bhowmik et al., 2017). While multivariate QMs attempt to address this issue (Piani and Haerter, 2012; Cannon, 2018), QM is inappropriate for the bias correction of future projections as it distorts future climate change in a nonphysical way (e.g., Maurer and Pierce, 2014). Other commonly used methods such as bias corrected spatial downscaling (BCSD, Wood et al., 2002), bias corrected constructed analogues (BCCA, Maurer et al., 2010), and multivariate adapted constructed analogues (MACA, Abatzoglou and Brown, 2012) also employ QM. Multivariate methods like localized constructed analogs (LOCA; Pierce et al., 2014, 2023) and asynchronous canonical correlation analysis (CCA; Bhowmik et al., 2017) offer improvements by better capturing the temporal covariability in climate variables (e.g., cross-correlation), but do not explicitly consider spatial dependence during bias correction. Finally, recent deep learning-based methods, like DeepSD (Vandal et al., 2017), leverage the power of deep learning algorithms to model complex, nonlinear relationships between climate variables and local physiographical information such as elevation, but have not been evaluated for spatiotemporal dependency in multiple climate variables. Additionally, none of the aforementioned methods, with the exception of LOCA v2 (Pierce et al., 2023), are geared towards reproducing observed extreme quantiles of climate variables. However, LOCA v2 also fails to preserve spatial dependency. We argue that bias correction should be viewed as a problem of density correction - calibrating the spatiotemporal joint density of the historical GCM projections to the spatiotemporal joint density of the observed multivariate climate data. Subsequently, the ‘calibration’ model will be run with future GCM projections to develop density-corrected (i.e., calibrated) GCM projections that remove the systematic biases in the GCM joint densities. Our effort here is to develop such a density correction approach by comparing the joint density fields of GCM projections with its observed counterpart.
@@
The proposed SPCDE connects conditional density estimation (CDE; Hyndman et al., 1996; Li and Racine, 2006) and bias correction in a unique manner, and in the process mitigates several drawbacks of commonly used bias correction methods. A key aspect of SPCDE (described in further detail in Sections 2.32.5) is the estimation of the joint distribution of GCM and observational data within a single SPQR model, and thereafter using cumulative density function (CDF) and quantile function (QF) transforms to calibrate the model data. Similar CDF based approaches have been used for bias correcting future climate projections (e.g., Michelangeli et al., 2009; Li et al., 2010); however, these approaches estimate the CDF of the GCM and observed data independent of each other. SPCDE also bears resemblance to normalizing flows (NFs; Rezende and Mohamed, 2015; Papamakarios et al., 2021; Kobyzev et al., 2021) and transport maps (Marzouk et al., 2016; Katzfuss and Schäfer, 2024). Both of these methods aim to transform a (usually complex) target distribution to a much simpler reference distribution (e.g., a standard Gaussian) through a series of invertible transformations; the transformations are referred to as flows and transport maps, respectively. Inverting the transformations allows for sampling from the target distribution as well as making predictions. Transport maps can be considered a special case of an NF (Katzfuss and Schäfer, 2024). Similarly, SPCDE parallels NFs - we can consider the GCM and observational data to be distributed according to the reference and target distributions respectively, with the flows being a sequence of CDFs and QFs. This also highlights the key difference between SPCDE and NFs. While the reference distribution in NFs is assumed to be known, the joint distribution of the GCM data also needs to be learned alongside the joint distribution of the observational data. While this is a more complex task, SPCDE’s ability to estimate them simultaneously makes the calibration step of density correction quite straightforward.
The remainder of this paper is organized as follows. Section 2 introduces the data and methods for our study. Section 3 contains a case study for density correction of climate model data at two different regions of the US. Section 4 concludes.
@@ 2.1 Overview
Section 2.2 describes the different datasets used in our study, including resolution and pre-processing used to align the different datasets on a common spatial grid. Section 2.3 introduces the SPCDE approach for density correction of TMAX and PRCP at a single grid cell. Sections 2.4 and 2.5 extend SPCDE to spatial and spatiotemporal settings, respectively. Section 2.6 outlines SPQR, the conditional density estimator that is used for SPCDE. Finally, Section 2.7 describes the metrics used to assess the quality of the calibrated data.
@@ 2.2 Temperature and precipitation datasets
2.3 Density correction model for GCM projections at a single location
Our proposed density correction model has three steps - estimation, projection, and calibration. We begin by stating our model for the general multivariate case in this section; explicit formulations for spatial and spatiotemporal data are described in subsequent sections. We build a statistical model for $p$ continuous responses, $\mbox{\bf Y}=(Y_{1},...,Y_{p})$ at a single location, which are assumed to have invertible marginal quantile and distribution functions that depends on variables Z (e.g., season or elevation). The model is described simultaneously for both climate model output and observational data, with an auxiliary variable $X=0$ indicating the observation is from the climate model and $X=1$ indicating it is from the observational record.
@@ 2.3 Density correction model for GCM projections at a single location
While (4) holds for any ordering of the $p$ variables, certain orderings will likely to lead to better estimation depending on the complexity of the $p$ densities. In the simple case of TMAX and PRCP at a single location (i.e., $p=2$), a natural ordering would be based on choosing $Y_{1}$ as TMAX and $Y_{2}$ as PRCP. This decomposes the conditional distribution of PRCP and TMAX as a conditional and a marginal distribution, assuming PRCP depends on TMAX. Physically, TMAX and PRCP are coupled, and choosing $Y_{1}$ as PRCP and $Y_{2}$ as TMAX also provides a valid statistical formulation. However, we assume that PRCP depends on TMAX for somewhat practical reasons. Intuitively, the distribution of TMAX has a much simpler functional form compared to PRCP. PRCP has a preponderance of zeros, and is therefore harder to estimate compared to TMAX. Our proposed dependence structure mitigates this by using TMAX as a predictor for PRCP. Our formulation is also validated in practice; SPCDE where TMAX depended on PRCP led to noticeably worse predictions over both the training and testing periods.
@@
Algorithm 1 Density correction of $Y_{1}=$ TMAX and $Y_{2}=$ PRCP at a single location
Data $\mbox{\bf Y},\mbox{\bf Z},X$
procedure SPQR for TMAX and PRCP
     Estimate $F_{1}$ and $Q_{1}$ for $Y_{1}|\mbox{\bf Z},X$triangleright Estimation step
     Estimate $F_{2}$ and $Q_{2}$ for $Y_{2}|Y_{1},\mbox{\bf Z},X$triangleright Estimation step
end procedure
procedure TMAX density correction
@@
procedure PRCP density correction
@@ 2.4 Extension to spatial data
Consider our problem of interest - density correction of TMAX and PRCP at multiple locations. If there are $n$ locations, a reasonable approach would be to first order them (e.g., based on their geographical coordinates), and then order TMAX and PRCP at each location. This could be represented by $\mbox{\bf Y}=(Y_{1},\ldots,Y_{p})$ for $p=2n$, and theoretically, Algorithm 1 from Section 2.3 can be applied. However, as the number of locations increase, so does the number of features $\mbox{\bf Y}_{(j)}$. This soon leads to computational problems, and directly applying (1) to large spatial (and spatiotemporal) data is computationally prohibitive. To mitigate this, we approximate (1) by means of a Vecchia approximation (Vecchia, 1988; Stein et al., 2004; Datta et al., 2016; Katzfuss and Guinness, 2021):
@@ 2.5 Extension to spatiotemporal data
Algorithm 2 Spatiotemporal density correction for TMAX and PRCP
@@
     procedure SPQR for TMAX and PRCP
         Estimate $F_{1l}$ and $Q_{1l}$ for $Y_{1lt}|\mbox{\bf Y}_{(1lt)},X$triangleright Estimation step
         Estimate $F_{2l}$ and $Q_{2l}$ for $Y_{2lt}|\mbox{\bf Y}_{(2lt)},X$triangleright Estimation step
     end procedure
     procedure TMAX density correction
@@
     procedure PRCP density correction
@@
In this study, we assume the conditional quantile function to be first order Markov in time, i.e., $T^{\prime}=1$. This assumption is reasonable given the selected are from the Southeast and Southwest US in regions where pronounced seasonality in precipitation is absent (Petersen et al., 2012). This implies that the Vecchia neighbor set for $Y_{1lt}$ is $\mbox{\bf Y}_{(1lt)}=\{Y_{1l,t-1},Y_{1,l-1,t},\ldots,Y_{11t}\}$, and for $Y_{2lt}$ is $\mbox{\bf Y}_{(2lt)}=\{Y_{2l,t-1},Y_{1l,t-1},Y_{1,l,t},\ldots,\\ Y_{1,l-m,t},Y_{2,l-1,t},\ldots,Y_{2,l-m,t}\}$. These are straightforward extensions of the Vecchia neighbor sets presented in Section 2.4, with TMAX on day $t$ being additionally depending on TMAX on day $t-1$, and PRCP on day $t$ additionally depending on both PRCP and TMAX from day $t-1$. The conditional QF $Q$, CDF $F$, and calibration function $C$, for $\mbox{\bf Y}_{lt}$ are defined identically to Section 2.3; these do not depend on $t$ due to the stationarity assumption. Algorithm 2 details the density correction methodology for spatiotemporal data.
@@ 2.6 Conditional density estimation using SPQR
$\Pi_{K}(\mbox{\bf X}):=\mbox{\bf x}^{(H+1)}$ $=\mbox{softmax}(\mbox{\bf W}^{(H+1)}\mbox{\bf x}^{(H)}+\mbox{\bf b}^{(H+1)}),$ (8)
@@
$\frac{\partial Q(\tau|\phi,\mbox{\bf X})}{\partial\tau}>0,\ \forall\ \mbox{\bf X},$
and does not require any second-stage monotonization treatment (Bondell et al., 2010).